OpenStudy (anonymous):

how many secs are in 1/2 hr

OpenStudy (2bornot2b):

\[0.5\times 60 \times 60 \]

OpenStudy (anonymous):

One hour is 60 minutes. One minute is 60 seconds. 60*60=360 seconds. One half of that is...

hero (hero):

60*60 is not 360